OpenStudy (anonymous):

if f(x) is twice differentiable defined for all x belongs to R such that f(x) =f(2-x) andf'(1/2)=f'(1/4) then min no. of values in[0,2] for which f''(x) vanishes is

OpenStudy (anonymous):

emm...what that means? "f''(x) vanishes"

OpenStudy (anonymous):

it means f''x=0

OpenStudy (anonymous):

f is symmetric with respect to x=1 and f''(x) vanishes once on the interval (0.25,0.5) so min number will b 2...but how to offer an acceptable reasoning

OpenStudy (anonymous):

\[f(x)=f(2-x)\]let x=1-x\[f(1-x)=f(1+x)\]f is symmetric with respect to x=1 so u just need to work on [0,1]

OpenStudy (anonymous):

and i believe using mean value theorem for the interval [0.25,0.5] will work here

OpenStudy (anonymous):

can we find the period of the function

OpenStudy (anonymous):

i think it is 2

OpenStudy (anonymous):

but the function is defined for all real values

OpenStudy (anonymous):

i m confused whether the function is periodic or not

OpenStudy (anonymous):

for all real values

OpenStudy (anonymous):

oh sorry its not periodic

OpenStudy (anonymous):

just with considering f(x) =f(2-x) we cant tell its periodic\[f(x)=(x-1)^2\]